# On the Influence of Pluto on Twotino Dynamics Through Their Mutual 4:3 Mean Motion Resonance

### v1

@@ 4.1 Comparison with Eris
One might think that Pluto affects the Twotinos because its aphelion reaches Twotinos’ semi-major axis. In this scenario, we would also expect Eris to affect the Twotinos and even the Plutinos since its orbit travels between 38 and 98 au; however, our studies have shown that Eris’s effect is at least an order of magnitude smaller than Pluto’s. This occurs even with Eris (plus Dysnomia) being $\sim 14\%$ more massive than Pluto (plus Charon).
When studying Twotino leaking rates, we found that, over 4 Gyr, considering the four giant planets, without Pluto, the fraction of stable Twotinos is close to 47%. If Eris is included as a massive perturber, the fraction of stable Twotinos decreases to 46%. Instead, when Pluto was included as a massive perturber (without Eris), this fraction fell to 19% (Muñoz-Gutiérrez et al., 2025). For Plutinos, the stable fractions correspond to 77%, 77%, and 36%, in the same previous scenarios, respectively. This is likely because, while neither Eris nor Pluto can trap TNOs in their MMRs, Neptune will trap TNOs in Pluto’s resonances, but not in Eris’s. Once locked in, Pluto can slightly yet significantly affect resonant objects on secular time-scales, while Eris will not.
@@ 5 Conclusions
We have shown that a single resonant argument of the 4:3 MMR between Pluto and Twotinos librates. This means that such a resonance is relevant for the secular evolution of Twotinos, and this could be the origin of their increased instability in simulations when Pluto is present, compared to cases where only the four giant planets are present. The amplitude libration of the 4:3 MMR is conventional, i.e., below $360^{\circ}$ for most Twotinos librating in the asymmetric islands of the 2:1 MMR with Neptune. For symmetric Twotinos this libration amplitude is much larger than the conventional $360^{\circ}$ typically considered in resonance studies, however, there is nonetheless a limited set of angular values that are preferentially visited in the co-rotating frame of Pluto; such behavior results in the same resonant phenomena as for Twotinos in the asymmetric islands, although with a weakened effect.
This result forces us to reconsider Pluto as an important ingredient when studying the long-term evolution of the trans-Neptunian region, especially for resonant populations such as Plutinos and Twotinos. With modern computers, its inclusion represents a minimal, almost imperceptible cost in integration time, so excluding it should not be considered a valid option.
@@ Acknowledgements.
This research was performed using services/resources provided by Grid UNAM, which is a collaborative effort driven by DGTIC and the research institutes of Astronomy, Nuclear Science and Atmosphere Science and Climate Change at UNAM. A.P.-V. acknowledges the DGAPA-PAPIIT grant IN112526. M.A.M.-G. acknowledges Universidad de Atacama for the DIUDA grant No. 88231R14.

### v2

@@ 4.1 Comparison with Eris
One might think that Pluto affects Twotinos because its aphelion reaches Twotinos’ semi-major axis. In this scenario, we would also expect Eris to affect Twotinos and even Plutinos since its orbit travels between 38 and 98 au; however, our studies have shown that Eris’s effect is at least an order of magnitude smaller than Pluto’s. This occurs even with Eris (plus Dysnomia) being $\sim 14\%$ more massive than Pluto (plus Charon).
In Paper I we studied Twotino leaking rates, and found that, over 4 Gyr, considering the four giant planets, without Pluto, the fraction of stable Twotinos is close to 47%. If Eris is included as a massive perturber, the fraction of stable Twotinos decreases to 46%. Instead, when Pluto was included as a massive perturber (without Eris), this fraction fell to 19%. For Plutinos, the stable fractions correspond to 77%, 77%, and 36%, in the same previous scenarios, respectively. This is likely because, while neither Eris nor Pluto can trap TNOs in their MMRs, Neptune will trap TNOs in Pluto’s resonances, but not in Eris’s. Once locked in, Pluto will slightly yet steadily affect resonant objects on secular time-scales, while Eris will not.
@@ 5 Conclusions
We have shown that a single resonant argument of the 4:3 MMR between Pluto and Twotinos librates. This means that such a resonance is relevant for the secular evolution of Twotinos, and this could be the origin of their increased instability in simulations when Pluto is present, compared to cases where only the four giant planets are included. The amplitude libration of the 4:3 MMR is conventional, i.e., below $360^{\circ}$ for most Twotinos librating in the asymmetric islands of the 2:1 MMR with Neptune. For symmetric Twotinos, the libration amplitude of up to $850^{\circ}$ exceeds the typically considered hard limit of $360^{\circ}$; nonetheless, there is a limited set of angular values that are preferentially visited by those Twotinos in the co-rotating frame of Pluto. Such behavior results in the same resonant phenomena as for Twotinos in the asymmetric islands, although with a weakened effect.
This result forces us to reconsider Pluto as an important component when studying the long-term evolution of the trans-Neptunian region, specifically for resonant populations such as Plutinos and Twotinos. With modern computers, the inclusion of Pluto represents a minimal effort, with almost imperceptible cost in integration time, so excluding it should not be considered a valid option.
@@ Acknowledgements.
We thank the anonymous referee for the careful review that helped us to improve this letter. This research was performed using services/resources provided by Grid UNAM, which is a collaborative effort driven by DGTIC and the research institutes of Astronomy, Nuclear Science and Atmosphere Science and Climate Change at UNAM. A.P.-V. acknowledges the DGAPA-PAPIIT grant IN112526. M.A.M.-G. acknowledges Universidad de Atacama for the DIUDA grant No. 88231R14.